An 900-g iron block is heated to 380 ∘C and placed in an insulated container (of negligible heat capacity) containing 36.0 g of water at 20.0 ∘C.

What is the equilibrium temperature of this system? The average specific heat of iron over this temperature range is 560 J/(kg⋅K).
Answer #1

Heat lost by iron = heat gained by water

let T be the final temperature

mironcironTiron = mwatercwaterTwater

0.9 * 560 * ( 380 - T) = 0.036 * 4186 *(T - 20)

191520 - 504 * T = 150.696T - 3013.92

194533.92 = 654.696 T

T = 297.14 oC

OR

T = 570.14 K